OpenStudy (anonymous):

how to solve use the formula h=40t - 5t^ to find the height in metes of a ball 5 second after it is thrown ??

OpenStudy (anonymous):

Use substuition. h=40*5-5*5^ h = 200-125 h = 75 That's the answer

OpenStudy (anonymous):

replace t by 5 to get \[40\times 5-5\times 5^2\] \[=200-5\times 25\] \[=200-125\] =\[75\]

OpenStudy (anonymous):

okay thank you very much :)